OpenStudy (anonymous):

Which of the following expressions is the conjugate of a complex number with 2 as the real part and 3i as the imaginary part? @primeralph

OpenStudy (anonymous):

anyone ?

OpenStudy (irishboy123):

for the conjugate of : \(z = 2 + 3i \) if \(z = a + i \ b\), then its conjugate \(\bar z\) or \(z^* = a - i \ b\)

OpenStudy (anonymous):

how do i set it up ?

OpenStudy (irishboy123):

you have \(z=2+3i \\\ = a+i b\) you now want \(\bar z\) or \(z^* = a-i b \)

OpenStudy (irishboy123):

so a = and b = and thus you have the conjugate

OpenStudy (anonymous):

im thinking its 3i + 2. am i right

OpenStudy (irishboy123):

go back a step if \(z=2+3i =a+i \ b\), what is a and what is b ... ?

OpenStudy (irishboy123):

which is \(z^∗=a−i \ b\)

OpenStudy (anonymous):

oh so 2+3i?

OpenStudy (irishboy123):

that's the number you started with you are looking for its conjugate

OpenStudy (anonymous):

totally confused

OpenStudy (irishboy123):

you started with \(z=2+3i\) you want its conjugate you get there by simply reversing the sign of the imaginary part